OpenStudy (anonymous):

PLEASE HELP!!!! MEDAL WILL BE REWARDED!!!! PLEASE HELP!!!! Find the first six terms of the sequence. a1 = -7, an = 2 • a(n-1) Answer choices: -7, -14, -12, -10, -8, -6 0, 2, -14, -12, -10, -8 -7, -14, -28, -56, -112, -224 -14, -28, -56, -112, -224, -448

sam (.sam.):

We have a1 = -7, -7 will be the starting number in the sequence

sam (.sam.):

Leaves you with -7, -14, -12, -10, -8, -6 or -7, -14, -28, -56, -112, -224

OpenStudy (anonymous):

I would say the third option?

sam (.sam.):

Then use their formula an = 2 • a(n-1) when a3=2(-7)(3-1)=-28

sam (.sam.):

Yes you got it

OpenStudy (anonymous):

Oh okay thank you so much for your help!!!!

sam (.sam.):

yw